OpenStudy (goformit100):

Find the number of different permutations of the letters of the word 'BANANA'.

OpenStudy (goformit100):

@shamim

OpenStudy (shamim):

ok. i m trying

OpenStudy (shamim):

there r 6 letters here

OpenStudy (shamim):

and there 3 A and 2 N here

OpenStudy (shamim):

so the number of permutation will b\[\frac{ 6! }{ 3!2! }\]

OpenStudy (shamim):

did u get it

OpenStudy (goformit100):

Ok then ?

OpenStudy (shamim):

\[\frac{ 6 \times 5 \times 4 \times 3! }{ 3! \times 2 }=60\]

OpenStudy (shamim):

just cancell 3!

OpenStudy (shamim):

feel free to ask any question